1.    It is claimed that the average weight of babies at birth is 3.4kg. The average weight of a random sample of 30 newly born babies was determined. It was found out that the average weight was 3.1kg. Is there a reason to believe that the average weight of babies at birth is not 3.4kg? Assume that the population standard deviation is 1.1kg. Use 0.05 level of significance.


1
Expert's answer
2021-04-14T14:10:18-0400

Ho: µ1 = µ0, (the average weight of babies is not different from 3.4kgs (µ0))


Ha: µ1 "\neq" µ0, (µ0 =3.4kg), (the average weight of babies is different from 3.4kgs)


Level of Significance: α=0.05


Test- statistic: Z- statistics (this is because the sample size is large enough, "n\geq 30" ). Thus we have "Z=\frac{\bar{X}-\mu_0}{\frac{s}{\sqrt{n}}}"


Tails in Distribution: Two-tailed


Reject H0 if "Z\geq 1.960 \text{ or if } Z\leq -1.960".


Test statistics


"Z=" "\frac{\bar{X}-\mu_0}{\frac{s}{\sqrt{n}}}=\frac{3.1-3.4}{\frac{1.1}{\sqrt{30}}}=-1.494"

We fail to reject H0 because "Z=-1.494>-1.960".


We do not have statistically significant evidence at α=0.05, to show that the average weight of babies at birth is not 3.4kg.